Determine the complex number z which satisfies the equations |z + 3i| = |z + 5 −2i|
and |z − 4i| = |z + 2i| simultaneously.
1
Expert's answer
2020-05-25T19:17:45-0400

Given, "\mid"z + 3i"\mid" = "\mid"z + 5 - 2i"\mid" ... eq. (1)

& "\mid"z - 4i"\mid" = "\mid"z + 2i"\mid" ... eq. (2)


Let z = x + iy.

Substituting the above value of z in equation (1) :


"\mid"x + iy + 3i"\mid" = "\mid"x + iy + 5 - 2i"\mid"

"\mid"x + i(y + 3)"\mid" = "\mid"(x + 5) + i(y - 2)"\mid"

"\sqrt{\smash[b]{ x² + (y + 3)²}}" = "\sqrt{\smash[b]{ (x + 5)² + (y - 2)²}}"


Squaring the above equation, we get,


x² + (y + 3)² = (x + 5)² + (y - 2)²

x² + y² + 9 + 6y = x² + 25 + 10x + y² + 4 - 4y

i.e., x² + y² + 9 + 6y - x² - 25 - 10x - y² - 4 + 4y = 0

y(6 + 4) - 10x + (9 - 25 - 4) = 0

10y - 10x - 20 = 0

i.e., 10y - 10x = 20 ... eq. (3)


Substituting z = x + iy in equation (2) :


"\mid"x + iy - 4i"\mid" = "\mid"x + iy + 2i"\mid"

"\mid"x + i(y - 4)"\mid" = "\mid"x + i(y+2)"\mid"

"\sqrt{\smash[b]{ x² + (y - 4)²}}" = "\sqrt{\smash[b]{ x² + (y + 2)²}}"


Squaring the above equation, we get,


x² + (y - 4)² = x² + (y + 2)²

x² + y² + 16 - 8y = x² + y² + 4 + 4y

i.e., x² + y² + 16 - 8y - x² - y² - 4 - 4y = 0

16 - 4 + y(-8 - 4) = 0

12 - 12y = 0

i.e., 12y = 12

"\therefore" y = 1


Substituting the above value of y = 1 in equation (3) :


10y - 10x = 20

10(1) - 10x = 20

10 - 10x = 20

i.e., 10x = 10 - 20 = - 10

"\therefore" x = -1


Therefore, z = -1 + i